OpenStudy (anonymous):

One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single?

OpenStudy (anonymous):

1. Take like terms together. x-0.6x = 384 0.4x=384 2. Find the value of x. x = 384/0.4 = 960 3. Take 60% of 960 60/100*960 = 576 Which is your answer.

OpenStudy (anonymous):

Because x = 960 people is the total number who download songs from that site every month, it is given that 60% of them download Q Sam's songs, so 60% of 960 would be 576.

OpenStudy (anonymous):

you are a life saver! thank you!(: